OpenStudy (anonymous):

The combined area of New York and California is 204,192 square miles. The area of California is 108,530 square miles more than that of New York. Find the land area of each state.

jhonyy9 (jhonyy9):

so let x area of New York and y area of California using this note how can you rewriting your exercise ?

jhonyy9 (jhonyy9):

x+y=204192 y=x+108530 so substitute this value of y inside first equation in place of y and will get the value of x

jhonyy9 (jhonyy9):

can you calcule it ?

jhonyy9 (jhonyy9):

x+x+108530=204192 2x=204192-108530 2x=95662 divide both sides by 2 x=95662/2 x=47831 so than you can calcule the value of y y=x+108530=47831+108530=156361 x=47831 y=156361

jhonyy9 (jhonyy9):

do you understand it ?

jhonyy9 (jhonyy9):

so than good luck bye

OpenStudy (anonymous):

Thank you that was easier than I thought .

jhonyy9 (jhonyy9):

yes i believe it good luck bye